# SemCSE-Multi: Multifaceted and Decodable Embeddings for Aspect-Specific and Interpretable Scientific Domain Mapping

### v1

@@ 5.3 Results
In contrast, both the Shuffled and Unconditioned control conditions yield markedly higher perplexity scores, thus clearly confirming that the decoder is not simply producing generic aspect-related text, but is instead leveraging the embedding to generate faithful reconstructions. Together, these results establish that embedding decoding provides a reliable and interpretable link between aspect-specific vector representations and their corresponding semantic descriptions. Exemplary reconstructions of summarizing sentences are included in Figure 3.
@@ 6 Interactive Embedding and Decoding
To obtain low-dimensional visualizations of the embedding space, we use the t-SNE method van der Maaten and Hinton (2008). The method begins by computing a pairwise distance matrix in the original high-dimensional space, which is converted into symmetric similarity probabilities $P$ that quantify how strongly two samples are related. It then optimizes low-dimensional coordinates $Y$ so that their corresponding similarity probabilities $Q$ mirror those from the high-dimensional space:
@@
This objective therefore ensures that points which are neighbors in the original embedding space remain close in the visualization, while unrelated points are pushed apart. For more details, see Appendix A.7.
@@ 6.2 Methods for Interaction
First, weighting different aspects is straightforward by adjusting the distance function used to compute the high-dimensional distance matrix $A$. For our multifaceted embeddings, we define a combined distance as $d(x_{i},x_{j})=\sum_{a}\alpha_{a}d_{a}(x_{i},x_{j})$, where $d_{a}(x_{i},x_{j})$ represents the distance between samples $i$ and $j$ with respect to aspect $a$, and the weights satisfy $\sum_{a}\alpha_{a}=1$. Modifying the values of $\alpha$ effectively prioritizes certain aspects over others, producing a low-dimensional visualization that reflects these relative importances. Figure 1 demonstrates the vastly differing visualizations that result from distinct aspect-specific weightings.
Second, embedding additional user-provided samples can be achieved by extending the matrix $A$ with the corresponding distances of the new sample in the high-dimensional space and optimizing the objective in Equation 1 with respect to just the one new low-dimensional embedding.
@@
Once obtained, this reconstructed embedding can be passed through the decoding pipeline (Section 5) to generate a natural language description, thereby revealing the semantic meaning associated with any location in the visualization.
For technical details and practical considerations regarding the optimization, see Appendix A.8.
@@ 6.3 Evaluation Setup
This is done by creating a t-SNE visualization for a single aspect (i.e., $\alpha_{a}=1$ and all other weights are 0) and mapping an additional sample represented by embedding $e_{n+1}$ into the visualization, yielding low-dimensional coordinates $y_{n+1}$. Afterwards, we remove that sample again and reconstruct a new high-dimensional embedding $\tilde{e}_{n+1}$ for the - now empty - position $y_{n+1}$. We then evaluate the perplexity of our decoder for embedding $\tilde{e}_{n+1}$, but with respect the the aspect-specific summaries for embedding $e_{n+1}$.
@@ 6.4 Results
The results of this evaluation are presented in Table 3. As expected, the reconstructed perplexity values are slightly higher than those obtained when decoding directly from the original (matching) high-dimensional embeddings. This increase is unsurprising, since the two-dimensional projection is not injective, so that multiple high-dimensional points could map to the same low-dimensional location.
Nevertheless, the perplexity remains only marginally higher than that of direct reconstruction, indicating that the reconstructed embeddings retain most of the semantic information. Most importantly, the perplexity is substantially lower than in the randomized control setup, demonstrating that the proposed approach effectively generates accurate natural language representations for arbitrary positions within the embedding visualization.
@@ 7 Conclusion
Based on the extensive experiments in the domain of invasion biology, complemented by evaluations in the medical domain (Appendix A.8), we believe that the improved capabilities for accurate and aspect-specific similarity assessments are applicable to a variety of domains, thus offering researchers powerful new tools to navigate and interpret large bodies of literature.
@@ A.1 Aspects
We identified six aspects of invasion biology that capture complementary dimensions of research studies that are relevant for assessing similarities between them. The aspects under consideration are the following:
@@ A.2 Summary Dataset Creation
We use Mistral Small 3.1 Mistral AI (2025) to generate four summarizing sentences for each sample in the dataset and for each aspect under consideration.
@@ A.3 Training Aspect Embedding Models
To train the individual aspect-specific embedding models, we use SciDeBERTa v2 Kim et al. (2023) as the base encoder (183M parameters). The 768-dimensional output embeddings are projected to a 150-dimensional space using an additional linear projection head. For each aspect, training is performed only on samples with at least two valid summaries (i.e., summaries other than “Not applicable.”). The corresponding dataset sizes are provided in Table 4.
@@ A.4 Training the Unified Model
For the unified model, we again use SciDeBERTa v2 Kim et al. (2023) as the base encoder (183M parameters), augmented with one linear projection head per aspect. Each projection head maps the 768-dimensional [CLS] token embedding to a 150-dimensional aspect-specific representation.
To construct the training and validation datasets, we use the same samples as before and embed all summaries for all aspects using the corresponding aspect-specific embedding models. For each sample and aspect, we then compute a single ground-truth aspect embedding by averaging the individual summary embeddings for that aspect. The unified embedding model is then trained to predict these aspect embeddings directly from the full scientific abstract, using the appropriate projection head.
@@ A.5 LLM Assessment Creation
A.6 Training Embedding Decoders
@@
To investigate the effect of model size and pre-training strategy, we evaluate two different LLMs: the purely autoregressive Llama 3 (8B) Grattafiori et al. (2024) and the 24B-parameter chat model Mistral Small 3.1 Mistral AI (2025).
@@
A.8 Details on Interactive Embedding and Decoding
@@
A.9 Details on Embedding Reconstruction Decoding

### v2

@@ 5.3 Results
In contrast, both the Shuffled and Unconditioned control conditions yield markedly higher perplexity scores, thus clearly confirming that the decoder assigns a higher likelihood to the ground-truth reconstruction if conditioned on the correct embedding. The BERTScore results show a similar picture, with BERTScores in the matching setting being consistently and substantially higher than those in the shuffled setting. This demonstrates that not only the likelihood of the ground truth is increased when conditioned on the correct embedding, but that the actual generations show a good alignment with the original ground-truth. Together, these results establish that embedding decoding provides a reliable and interpretable link between aspect-specific vector representations and their corresponding semantic descriptions. Exemplary reconstructions of summarizing sentences are included in Figure 3.
@@ 6 Interactive Embedding and Decoding
To obtain low-dimensional visualizations of the embedding space, we use the t-SNE method van der Maaten and Hinton (2008). The method begins by computing a pairwise distance matrix $A$ in the original high-dimensional space, which is converted into symmetric similarity probabilities $P$ that quantify how strongly two samples are related. It then optimizes low-dimensional coordinates $Y$ so that their corresponding similarity probabilities $Q$ mirror those from the high-dimensional space:
@@
This objective ensures that points which are neighbors in the original embedding space remain close in the visualization, while unrelated points are pushed apart. For details, see Appendix A.9.
@@ 6.2 Methods for Interaction
First, weighting different aspects is straightforward by adjusting the distance function used to compute the distance matrix $A$. Using individual, aspect-specific distances $d_{a}(x_{i},x_{j})$ between samples $i$ and $j$ for aspect $a$, we define a combined distance as $d(x_{i},x_{j})=\sum_{a}\alpha_{a}d_{a}(x_{i},x_{j})$ with $\sum_{a}\alpha_{a}=1$, so that modifying the values of $\alpha$ effectively prioritizes certain aspects over others, producing a low-dimensional visualization that reflects these relative importances. Figure 1 demonstrates the vastly differing visualizations that result from distinct aspect-specific weightings.
Second, embedding additional user-provided samples is done by extending the matrix $A$ with the distances of the new sample in the embedding space and optimizing Equation 1 with respect to just the new low-dimensional coordinates.
@@
Once obtained, this reconstructed embedding can be passed through the decoding pipeline to generate a natural language description, thereby revealing the semantic meaning associated with any location in the visualization. For technical details and practical considerations regarding the optimization, see Appendix A.10.
@@ 6.3 Evaluation Setup
This is done by creating a t-SNE visualization for a single aspect (i.e., $\alpha_{a}=1$ and all other weights are 0) and mapping an additional sample represented by embedding $e_{n+1}$ into the visualization, yielding low-dimensional coordinates $y_{n+1}$. Afterwards, we remove that sample again and reconstruct a new high-dimensional embedding $\tilde{e}_{n+1}$ for the (now empty) position $y_{n+1}$. We then evaluate our decoder on embedding $\tilde{e}_{n+1}$, but taking aspect-specific summaries for embedding $e_{n+1}$ as targets.
@@ 6.4 Results
Evaluation results are presented in Table 3. As expected, the reconstructed scores are slightly underperforming those obtained when decoding directly from the original (matching) high-dimensional embeddings. This is unsurprising since the two-dimensional projection is not injective, so that multiple high-dimensional points could map to the same low-dimensional location. Nevertheless, both scores are approaching those of direct reconstruction and, most importantly, substantially outperform the shuffled control setup, thus demonstrating that the proposed approach effectively generates accurate natural language representations for arbitrary positions within the embedding visualization.
@@ 7 Conclusion
Based on the extensive experiments in the domain of invasion biology, complemented by evaluations in the medical domain (Appendix B), we believe that the improved capabilities for accurate and aspect-specific similarity assessments are applicable to a variety of domains, thus offering researchers powerful new tools to navigate and interpret large bodies of literature.
@@ A.1 Aspects
With the help of domain experts, we identified six aspects of invasion biology that capture complementary dimensions of research studies that are relevant for assessing similarities between them. These aspects were chosen based on two main properties: the usefulness of them for domain experts in the context of visualization and literature search, as well as the prevalence of those aspects within scientific abstracts. The latter point is crucial, since, while other aspects that only occur in the full-text might be important, this information is not available in our training pipeline, and would also not be available in common downstream settings like literature search. The aspects under consideration are the following:
@@ A.2 Summary Dataset Creation
We use Mistral Small 3.1 Mistral AI (2025) to generate four summarizing sentences for each sample in the dataset and for each aspect under consideration. This model in particular was selected due to good performance in other (unpublished) experiments in the domain of invasion biology, thus making it the best candidate of comparable size at the time of summary generation.
@@ A.3 Training Aspect Embedding Models
To train the individual aspect-specific embedding models, we use SciDeBERTa Kim et al. (2023) as the base encoder (183M parameters), since it has been used by the original SemCSE paper and displayed strong results. Also using the same model eliminates a variable that could cause differences in results, so that using the same model enables a more precise relative performance assessment. The 768-dimensional output embeddings are projected to a 150-dimensional space using an additional linear projection head. For each aspect, training is performed only on samples with at least two valid summaries (i.e., summaries other than “Not applicable.”). The corresponding dataset sizes are provided in Table 4.
@@ A.4 Training the Unified Model
For the unified model, we again use SciDeBERTa Kim et al. (2023) as the base encoder (183M parameters), augmented with one linear projection head per aspect. Each projection head maps the 768-dimensional [CLS] token embedding to a 150-dimensional aspect-specific representation.
To construct the training and validation datasets, we use the same samples as before and embed all summaries for all aspects using the corresponding aspect-specific embedding models. For each sample and aspect, we then compute a single ground-truth aspect embedding by averaging the individual summary embeddings for that aspect. The unified embedding model is then trained to predict these aspect embeddings directly from the full scientific abstract, using the appropriate projection head. To do this, we leverage an L2 loss, since it is the natural choice to facilitate precise reconstruction of known targets.
@@ A.5 LLM Assessment Creation
A.6 Statistical Significance of Correlations
We evaluate the statistical significance of having higher correlation scores on the diagonals (i.e., between matching embedding and ground-truth aspects) compared to off-diagonal cells (i.e., cells that evaluate embeddings for one aspects against ground-truth for another). To this end, we perform a permutation-based significance test by creating the matrix $S_{a}\in\mathbb{R}^{n\times n}$ that contains the similarity scores between all pairs of $n$ scientific abstracts calculated using embeddings for aspect $a$, and the matrix $S_{b}\in\mathbb{R}^{n\times n}$ containing the same scores for embeddings of aspect $b$. We then repeatedly create a randomized matrix $S_{r}$ that contains, in each cell, either the score in the same cell of matrix $S_{a}$ or the score in the same cell of matrix $S_{b}$. We then evaluate the percentage of matrices $S_{r}$ that display a correlation with the ground truth of aspect $a$ that is at least as high as that for matrix $S_{a}$. For each possible pair of aspects and within 10,000 permutations, none of the permuted matrices showed a correlation as high as in the matching setting, thus demonstrating statistical significance. Exceptions are the (generally under-performing) recommendation aspect, as well as the case with hypothesis embeddings being evaluated against research question ground-truth, since the hypothesis embeddings already show stronger alignment than the matching embeddings even without randomization.
A.7 A Unified Baseline
SemCSE-Multi is based on a two-stage training procedure that first trains individual, aspect-specific embedding models using aspect-summaries and subsequently distills a unified model that predicts these embeddings directly from the abstract.
In this section, we evaluate an additional baseline that directly trains a unified model for all aspects at once in a single step. This is done by using the model structure of SemCSE-Multi (i.e., base encoder with one linear projection head for each aspect) and using the contrastive objective from Section 3. For each parameter update, the contrastive objective is performed once for a separate batch for each aspect (and thus only applied to the specific linear projection head), thus updating the specific part of the model to gain a more precise understanding of that aspect.
Further, we now include complete abstracts in the objective, since otherwise the model would not learn to predict accurate embeddings from these longer and more information-rich texts. The task is therefore to match summaries of the same aspect to each other, and also to match abstracts to their aspect-specific summaries.
We find that the correlations between embedding-similarities and LLM-scores for matching aspects (i.e., the five on-diagonal scores, excluding the underperforming recommendation aspect) are similar (42.6 for SemCSE-Multi and 42.8 for the baseline). However, the eighteen off-diagonal elements (excluding the cells matching hypothesis to research question, since these are sensible to be quite high) show a different picture: here, SemCSE-Multi achieves an average score of 21.43 compared to 28.3 for the baseline.
The reason for the reduced ability to disentangle the individual aspects is likely the joined process of training the baseline model for all aspects as once. While each aspect has its own embedding, these embeddings are all created by applying a linear projection to the same, shared, high-dimensional embedding produced by the encoder. For this reason, the update made to this embedding by one aspect can also affect the embedding of another aspect, and as long as this update does not reduce the matching performance on this other aspect, this change is not counteracted at any point. This ultimately leads to shared information between different aspects, and thus to reduced disentanglement.
Training each aspect in isolation instead ensures that only information useful for that aspect is included in the embedding, instead of also having irrelevant information from other aspects included.
A.8 Training Embedding Decoders
@@
To investigate the effect of model size and pre-training strategy, we evaluate two different LLMs: the purely autoregressive Llama 3 (8B) Grattafiori et al. (2024) and the 24B-parameter chat model Mistral Small 3.1 Mistral AI (2025). These models in particular were chosen due to, for the case of Mistral, being the model that generated the summaries in the first place, and, in the case on Llama, being simply a popular and smaller autoregressive LLM that has also been used in the original SemCSE paper Brinner and Zarriess (2025).
@@
A.10 Details on Interactive Embedding and Decoding
@@
A.11 Details on Embedding Reconstruction Decoding